OpenStudy (anonymous):

Solve b^2+15b+54=0 what is the solution?

OpenStudy (anonymous):

(b+6)(b+9)

hero (hero):

b^2 + 9x+ 6x = 54 b(x+9)+6(x+9) (x+9)(x+6)=0 b = -9,-6 you have to find b values

hero (hero):

okay, my solution is a little jacked up

OpenStudy (anonymous):

so i would put tht is a solution form

hero (hero):

Yeah, but let me fix it

hero (hero):

b^2 + 9b+ 6b + 54 b(b+9)+6b+9) (b+9)(b+6)=0 b = -9,-6

OpenStudy (anonymous):

ty